OpenStudy (anonymous):

Reduce the following rational expression to lowest terms, if possible. (Problem in comments)

OpenStudy (anonymous):

\[\frac{ 4x^2+x-14 }{ x^4-16 }\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

lol, thank you

OpenStudy (anonymous):

@Jack1 can you help my friend with her work as well please ?

OpenStudy (jack1):

\[\large \frac {4x^2+x−14}{x^4−16}\] 2 ways to start, either flip the signs on the powers \[\large 4x^2+x−14 \times (x^{\color{red}{-4}}−16^{\color{red}{-1}})\], then expand through multiplication or (easier) factorise the top line first (numerator): \[\large \frac {4x^2+x−14}{x^4−16}\]\[\large \frac {(4x-7)(x+2)}{x^4−16}\] then factorise the bottom line (denominator): \[\large \frac {(4x-7)(x+2)}{(x^2−4)(x^2 + 4)}\] \[\large \frac {(4x-7)(x+2)}{(x−2)(x+2)(x^2 + 4)}\] now we can cancel out the factors (x+2)

OpenStudy (anonymous):

\[\frac{ (4x-7) }{ (x-2)(x^2+4) }\]

OpenStudy (jack1):

perfect. you can re-expand the bottom line to a single equation if u like, but otherwise that's pretty mych it

OpenStudy (anonymous):

that's the answer? can i input it like this?

OpenStudy (jack1):

yup, or \[\large \frac {4x-7}{x^3-2 x^2+4 x-8}\], whichever, they're both correct

OpenStudy (anonymous):

thank you so much !!!

OpenStudy (jack1):

neh, all goods, welcomes! anything for a friend of @Shay17 's ;) slaters

OpenStudy (anonymous):

where are you from?

OpenStudy (anonymous):

Lol jack is the best! Thanks for the love ;)

OpenStudy (anonymous):

^_^

OpenStudy (jack1):

;)

OpenStudy (anonymous):

where are you from @Jack1 ?

OpenStudy (jack1):

Gold Coast, QLD (Down Unda) U @cookiibabii93 ?

OpenStudy (anonymous):

Austrailia lol (Down Unda) that's a first

OpenStudy (anonymous):

omg, thats sooo aweosme, i could never go to australia tho lol too many scary creatures lol im from america, califronia. excuse the bad grammer, i just dont feel like capitaizing and punctuating words right now lol

OpenStudy (anonymous):

you never heard anyone say down unda, shay? lol

OpenStudy (jack1):

noice catch Shay, I see u sheila's caught onto the accent there (unda) ;D

OpenStudy (anonymous):

Nope lol

OpenStudy (anonymous):

Sheila??? 0_o

OpenStudy (anonymous):

lmaoo they say it all the time on the outback steak house commercials!! lmaoo

OpenStudy (anonymous):

sheila is like saying females lol

OpenStudy (anonymous):

omg, shay lol

OpenStudy (anonymous):

Ohh see i dnt know this stuff

OpenStudy (jack1):

nah, they're fine hey, though last time i was up north we kinda had to abandon the beach as a 4m croc decided to go for a dip also ;) other than that, the creatures are pretty ... i don't know, standard ?

OpenStudy (jack1):

yup, sheila's = girls

OpenStudy (anonymous):

lol i would've had a heart attack!

OpenStudy (anonymous):

lol oh no! i would of ran all the way back to america lol

OpenStudy (anonymous):

he was in america lol

OpenStudy (anonymous):

up north right ?

OpenStudy (anonymous):

so how do u say males? mates or is that just a general word for friend?

OpenStudy (anonymous):

i think he meant north australia

OpenStudy (jack1):

the lifeguards caught it well before we did, he drifted by after a few hours tho and everything went back to normal, down here at the Goldy we don't get anything like that tho, the occasional shark, but nothin crazy

OpenStudy (anonymous):

ohh ok

OpenStudy (jack1):

ah, yep, north queensland, went up to Rockhampton for a bit

OpenStudy (jack1):

males = blokes, mates are mates (not specific gender)

OpenStudy (anonymous):

where is rockhampton ?

OpenStudy (anonymous):

i hesitate & cross the street when i see a dog!! lmaoo

OpenStudy (anonymous):

i only thought people from england & britian said blokes.

OpenStudy (jack1):

Rocky is...i dunno, about 8 hrs drive away, so prolly 800 kms, give or take, its just inside the Tropic of Capricorn, ann humid as in summer hey

OpenStudy (jack1):

England's full of Aussies, Kiwis and South Africans, pretty sure most of the ex-colonies picked "blokes" up from them or them from us... dunno?

OpenStudy (anonymous):

lol, you are sooo awesome!

OpenStudy (anonymous):

how old are you?

OpenStudy (jack1):

almost 17, u girls?

OpenStudy (anonymous):

wow i'm 17 didn't think you were that young

OpenStudy (anonymous):

18 in july

OpenStudy (anonymous):

i just turned 21...you are sooo smart! lol me neither

OpenStudy (jack1):

im totes mature though ;D

OpenStudy (anonymous):

I see lol

OpenStudy (anonymous):

what's totes ? lol

OpenStudy (jack1):

nah, my family are study freaks, one sisters a vet, the others already a doctor so can't fall behind, we're kinda competative ;D

OpenStudy (anonymous):

totes=very?

OpenStudy (jack1):

totes : totally ;D

OpenStudy (anonymous):

ohh okay ^_^

OpenStudy (anonymous):

lol ima start using totes lol

OpenStudy (anonymous):

I wish i wasnt from America lol the damn education system, suuuucks dude!! lol

OpenStudy (jack1):

yeah i heard it's a bit rough over there, but your uni's are top notch tho!

OpenStudy (anonymous):

what's uni's ? lol

OpenStudy (jack1):

universities... ummm u guys call em collages or somethin, yeah?

OpenStudy (anonymous):

lmaoo shay! i was gonna ask the same thing though

OpenStudy (anonymous):

ohh yea college lol

OpenStudy (anonymous):

universities too i just never heard of uni's

OpenStudy (anonymous):

lol f college! lol pellet is expensive as hell !! ugh lol

OpenStudy (anonymous):

omg, not pelletes, S H I T's

OpenStudy (anonymous):

@cookiibabii93 You can't curse on here please be mindful ! Lol

OpenStudy (anonymous):

Don't curse and college is good for you ! Lol

OpenStudy (anonymous):

lmaooo, ugggh !

OpenStudy (jack1):

rully, when we finish school, uni's paid for by the govt so long as ur grades are good enough for the course u want, but it's more like a load, once u graduate u have to pay it all back.... k night pretty ladies, was awesome chattin, imma dinner now then study, c u online another night?

OpenStudy (anonymous):

Okay and it's morning in here so goodnight to you and goodmorning to us

OpenStudy (anonymous):

I want some dinner lol goodnight! i'd call you handsome, but i dont want to be a pedophile or something lol

OpenStudy (jack1):

hahaha, i think it's only underage girls that get ppl in trouble, it's never unwelcome attention coming FROM girls, just look at all the cougars at One Direction concerts xD

OpenStudy (jack1):

slaters and g'night!

OpenStudy (anonymous):

lmaoo, what's slaters?

OpenStudy (anonymous):

i asked that too b4 he said that to me

OpenStudy (anonymous):

i forgot tho i think it means later he said

OpenStudy (anonymous):

awesomeeeeeeeee!!!